OpenStudy (anonymous):

the area of a parking lot is 2475 ft^2. The rectangular parking lot has dimensions such that the length is 10 feet longer than the width. What is the dimensions of the praking lot?

OpenStudy (anonymous):

l=10+w and area = l*w = (10+w)*w w^2+10w-2475=0

OpenStudy (anonymous):

\[w=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] b= 10, c=-2475, a=1 plug in the values